# Green AI: A Preliminary Empirical Study on Energy Consumption in DL Models Across Different Runtime Infrastructures

### v1

@@ Abstract.
Deep Learning (DL) frameworks such as PyTorch and TensorFlow include runtime infrastructures responsible for executing trained models on target hardware, managing memory, data transfers, and multi-accelerator execution, if applicable. Additionally, it is a common practice to deploy pre-trained models on environments distinct from their native development settings. This led to the introduction of interchange formats such as ONNX, which includes its runtime infrastructure, and ONNX Runtime, which work as standard formats that can be used across diverse DL frameworks and languages. Even though these runtime infrastructures have a great impact on inference performance, no previous paper has investigated their energy efficiency. In this study, we monitor the energy consumption and inference time in the runtime infrastructures of three well-known DL frameworks as well as ONNX, using three various DL models. To have nuance in our investigation, we also examine the impact of using different execution providers. We find out that the performance and energy efficiency of DL are difficult to predict. One framework, MXNet, outperforms both PyTorch and TensorFlow for the computer vision models using batch size 1, due to efficient GPU usage and thus low CPU usage. However, batch size 64 makes PyTorch and MXNet practically indistinguishable, while TensorFlow is outperformed consistently. For BERT, PyTorch exhibits the best performance. Converting the models to ONNX usually yields significant performance improvements but the ONNX converted ResNet model with batch size 64 consumes approximately 10% more energy and time than the original PyTorch model.
@@ 1. Introduction
We find out that, on one hand, inference time and energy consumption exhibit a very strong correlation (spearman’s rho=0.99, p << 0.0001) when considering the aggregated results of the study. On the other hand, the performance and energy efficiency of DL are fickle and difficult to predict. MXNet outperforms both PyTorch and TensorFlow for vision models in batch size 1, due to efficient GPU usage and thus low CPU usage. However, batch size 64 makes PyTorch and MXNet practically indistinguishable, while TensorFlow is outperformed consistently. For BERT, PyTorch outperforms the other frameworks and MXNet is surprisingly inefficient with batch size 1. Converting the models to ONNX usually yields significant improvements in energy efficiency and inference time. Nevertheless, the ONNX-converted ResNet model with batch size 64 consumes approximately 10% more energy and time than the original PyTorch model. Finally, in our preliminary investigation of execution providers, we observe that TensorRT always outperforms CUDA.
@@ 2. Related Work
Openja et al. (Openja et al., 2022) presented an empirical study assessing the effectiveness of conversion formats, ONNX and CoreML, in the accuracy, performance, and robustness of pre-trained DL models. The study used Keras and PyTorch to train five widely used models across three datasets. Results indicate that the accuracy is the same while converted models have smaller sizes. Similarly, in a very recent study, Jacques et al. (Jacques et al., 2024) compared the inference time and battery usage of DL models and their converted version on iOS devices. Georgiou et al. (Georgiou et al., 2022) conducted experiments on PyTorch and TensorFlow in both training and inference, aiming to gauge the energy efficiency and performance of each one. They further investigated the trade-off between resource usage and accuracy. While the methodologies of the three recent studies closely align with ours, It is worth mentioning that our investigation encompasses a different scope. We include MXNet in addition to TensorFlow and PyTorch, and for the ONNX converted models, we delve deeper by examining its performance on two different execution providers: CUDA and TensorRT. Furthermore, it’s worth highlighting that energy considerations were not accounted for in the first study (Openja et al., 2022).
@@ 3. Methodology
How do the hardware accelerator libraries affect the performance of inference for models running on ONNX?
@@ 4.1. Energy efficiency of different frameworks
We proceed with our investigation in a similar manner as in RQ1 and further examine batch sizes of 1 and 64, this time focusing on exporting the ONNX format of each pre-trained model and running the inference of the exported model on the ONNX Runtime. Table  1 shows that the previously observed poor performance of TensorFlow at batch size 1 is mitigated after conversion, highlighting an optimization improvement achieved through ONNX. In batch size of 1, beyond TensorFlow, ONNX models not only outperform the original models by demonstrating an improvement in inference time across all cases but also contribute to a reduction in energy consumption. With a batch size of 64, in the ONNX versions converted from MXNet in both MobileNet and ResNet models, there is an observed increase of around 13% in inference time and approximately 10% increase in total energy consumption. Similarly, the PyTorch to ONNX for MobileNet model exhibits an increase of approximately 10% in both energy consumption and inference time. These results further reinforce the point made for RQ1: if inference time and energy efficiency are critical, experimentation is required because there is no overall winning solution, even when employing a runtime aimed at making inference efficient.

### v2

@@ Abstract.
Deep Learning (DL) frameworks such as PyTorch and TensorFlow include runtime infrastructures responsible for executing trained models on target hardware, managing memory, data transfers, and multi-accelerator execution, if applicable. Additionally, it is a common practice to deploy pre-trained models on environments distinct from their native development settings. This led to the introduction of interchange formats such as ONNX, which includes its runtime infrastructure, and ONNX Runtime, which work as standard formats that can be used across diverse DL frameworks and languages. Even though these runtime infrastructures have a great impact on inference performance, no previous paper has investigated their energy efficiency. In this study, we monitor the energy consumption and inference time in the runtime infrastructures of three well-known DL frameworks as well as ONNX, using three various DL models. To have nuance in our investigation, we also examine the impact of using different execution providers. We find out that the performance and energy efficiency of DL are difficult to predict. One framework, MXNet, outperforms both PyTorch and TensorFlow for the computer vision models using batch size 1, due to efficient GPU usage and thus low CPU usage. However, batch size 64 makes PyTorch and MXNet practically indistinguishable, while TensorFlow is outperformed consistently. For BERT, PyTorch exhibits the best performance. Converting the models to ONNX yields significant performance improvements in the majority of cases. Finally, in our preliminary investigation of execution providers, we observe that TensorRT always outperforms CUDA.
@@ 1. Introduction
We find out that, on one hand, inference time and energy consumption exhibit a very strong correlation (spearman’s rho=0.99, p <<< 0.0001) when considering the aggregated results of the study. On the other hand, the performance and energy efficiency of DL are fickle and difficult to predict. MXNet outperforms both PyTorch and TensorFlow for vision models in batch size 1, due to efficient GPU usage and thus low CPU usage. However, batch size 64 makes PyTorch and MXNet practically indistinguishable, while TensorFlow is outperformed consistently. For BERT, PyTorch outperforms the other frameworks and MXNet is surprisingly inefficient with batch size 1. Converting the models to ONNX usually yields significant performance improvements with a few exceptions for ResNet and Mobilenet with batch size 64. In these cases, converted models of MXNet and PyTorch consume around 10-13% more energy and time than the original models. Finally, in our preliminary investigation of execution providers, TensorRT always outperforms CUDA.
@@ 2. Related Work
Openja et al. (Openja et al., 2022) presented an empirical study assessing the effectiveness of conversion formats, ONNX and CoreML, in the accuracy, performance, and robustness of pre-trained DL models. The study used Keras and PyTorch to train five widely used models across three datasets. Results indicate that the accuracy is the same while converted models have smaller sizes. Similarly, in a very recent study, Jacques et al. (Jacques et al., 2024) compared the inference time and battery usage of DL models and their converted versions on iOS devices. Georgiou et al. (Georgiou et al., 2022) conducted experiments on PyTorch and TensorFlow in both training and inference, aiming to gauge the energy efficiency and performance of each one. They further investigated the trade-off between resource usage and accuracy. While the methodologies of the three recent studies closely align with ours, It is worth mentioning that our investigation encompasses a different scope. We include MXNet in addition to TensorFlow and PyTorch, and for the ONNX converted models, we delve deeper by examining its performance on two different execution providers: CUDA and TensorRT. Furthermore, it’s worth highlighting that energy considerations were not accounted for in the first study (Openja et al., 2022).
@@ 3. Methodology
How do the execution providers affect the performance of inference for models running on ONNX?
@@ 4.1. Energy efficiency of different frameworks
We proceed with our investigation in a similar manner as in RQ1 and further examine batch sizes of 1 and 64, this time focusing on exporting the ONNX format of each pre-trained model and running the inference of the exported model on the ONNX Runtime. Table  1 shows that the previously observed poor performance of TensorFlow at batch size 1 is mitigated after conversion, highlighting an optimization improvement achieved through ONNX. In batch size of 1, beyond TensorFlow, ONNX models not only outperform the original models by demonstrating an improvement in inference time across all cases but also contribute to a reduction in energy consumption. With a batch size of 64, in the ONNX versions converted from MXNet in both MobileNet and ResNet models, there is an observed around 13% increase in inference time and approximately 10% increase in total energy consumption. Similarly, the ONXX MobileNet model converted from PyTorch exhibits an increase of approximately 10% in both energy consumption and inference time. These results further reinforce the point made for RQ1: if inference time and energy efficiency are critical, experimentation is required because there is no overall winning solution, even when employing a runtime aimed at making inference efficient.